A 52.0 mL aliquot of a 1.20M solution is diluted to a total volume of 208mL. A 104 mL portion of that solution is diluted by adding 139mL of water. what is the final concentration? assume the volumes are additive.
1
Expert's answer
2019-09-25T08:21:18-0400

As number of moles of solute remains constant throughout the dilution;

Moles in 52 ml solution"=52 \times 1.2 \times 10^{-3}=52 \times 1.2\space mili mol" which is equal to the number of moles in 208 ml solution.

Moles in 104 solution"=\frac{1}{2}\times 52 \times 1.2 \times 10^{-3} moles" "=26 \times 1.2 \times 10^{-3} moles"

Final volume "=104+139=243 ml"

Final concentration"=\frac{ 26 \times 1.2 \times 10^{-3}}{243 \times 10^{-3}}=0.128M"